OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 49x2 − 28x + 16 4a2 − 20a + 25 25b2 − 20b − 16 16x2 − 24x − 9

OpenStudy (anonymous):

pls explain

OpenStudy (skullpatrol):

The one with this form: $$a^2 +2ab+b^2$$

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

$$a^2+2ab+b^2=(a+b)^2$$

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

Let's look at the first choice: $$49x^2 − 28x + 16 =(7x)^2 -2*14x+4^2$$ Does that fit the "perfect square" pattern?

OpenStudy (anonymous):

no?

OpenStudy (skullpatrol):

Ok, can you check the others?

OpenStudy (anonymous):

i guess i can try

OpenStudy (anonymous):

i dont really get how to

OpenStudy (skullpatrol):

Write each one in this form first$$4a^2 − 20a + 25=(2a)^2-2*10a+5^2$$

OpenStudy (anonymous):

i still dint get it

OpenStudy (skullpatrol):

Perfect squares look like this $$a^2+2ab+b^2=(a+b)^2$$ $$a^2-2ab+b^2=(a-b)^2$$

OpenStudy (anonymous):

what happenned to the 2

OpenStudy (skullpatrol):

So the first two do not fit the pattern, right?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

What do you mean "what happenned to the 2" ???

OpenStudy (anonymous):

in 2ab whear did it go

OpenStudy (skullpatrol):

Where "in 2ab"?

OpenStudy (skullpatrol):

Here is the 3rd one: $$ 25b^2 − 20b − 16=(5b)^2-2*10b-4^2$$ does it fi the pattern?

OpenStudy (skullpatrol):